# Hai-Long Sun1,2,3   Da-Wei Zhou1,2  Yang Li3  Shiyin Lu3  Chao Yi1,2 Qing-Guo Chen3  Zhao Xu3  Weihua Luo3  Kaifu Zhang3  De-Chuan Zhan1,2  Han-Jia Ye1,2

### v2

@@ Abstract
The rapid development of Multimodal Large Language Models (MLLMs) like GPT-4V has marked a significant step towards artificial general intelligence. Existing methods mainly focus on aligning vision encoders with LLMs through supervised fine-tuning (SFT) to endow LLMs with multimodal abilities, making MLLMs’ inherent ability to react to multiple languages progressively deteriorate as the training process evolves. We empirically find that the imbalanced SFT datasets, primarily composed of English-centric image-text pairs, lead to significantly reduced performance in non-English languages. This is due to the failure of aligning the vision encoder and LLM with multilingual tokens during the SFT process. In this paper, we introduce Parrot, a novel method that utilizes textual guidance to drive visual token alignment at the language level. Parrot makes the visual tokens condition on diverse language inputs and uses Mixture-of-Experts (MoE) to promote the alignment of multilingual tokens. Specifically, to enhance non-English visual tokens alignment, we compute the cross-attention using the initial visual features and textual embeddings, the result of which is then fed into the MoE router to select the most relevant experts. The selected experts subsequently convert the initial visual tokens into language-specific visual tokens. Moreover, considering the current lack of benchmarks for evaluating multilingual capabilities within the field, we collect and make available a Massive Multilingual Multimodal Benchmark which includes 6 languages, 15 categories, and 12,000 questions, named as MMMB. Our method not only demonstrates state-of-the-art performance on multilingual MMBench and MMMB, but also excels across a broad range of multimodal tasks. Code is available at: https://github.com/AIDC-AI/Parrot.
@@ 1 Introduction
The rapid development of Large Language Models (LLMs), such as GPT-4 [52; 8; 45; 46; 53; 59; 66], has gained significant attention. However, LLMs are limited to processing a single textual modality. The expansion into visual modalities has endowed LLMs with multimodal capabilities [64; 2; 63; 75; 17; 27; 32; 65; 72; 7], thereby accelerating the development of Multimodal Large Language Models (MLLMs) and further bringing us closer to the realization of Artificial General Intelligence (AGI).
Current MLLMs mainly rely on pre-trained LLMs and vision encoders, focusing on bridging the modality gap by aligning visual features with language embedding tokens. Existing research predominantly employs either a Q-Former [31; 17; 6] or an MLP projector [38; 35; 12; 34; 37; 42] to align vision encoders with LLMs. The training data [38] mainly consists of English-centric data from image captions and multimodal conversations. During training, the alignment component converts the visual features into language embedding tokens. With encoded visual features, the LLM can integrate visual information to respond to multimodal inputs.
Multilingual capability in MLLMs entails the ability to generate responses in the same language as the input, catering to the linguistic diversity inherent in conversation. Therefore, it is critically important for processing language-specific content and cultural differences, ensuring equitable access to technological benefits for individuals across diverse regions and nations [13; 24]. Many LLMs possess multilingual capabilities [60; 5; 46], enabling diverse language responses according to user input. However, after the alignment training of MLLMs, the model may lose its ability to understand, process, or generate in non-English languages, and we call this phenomenon multilingual erosion. For example, LLaVA [38] usually responds in English, regardless of the input language. Therefore, it is essential to enhance MLLM’s multilingual capabilities during multimodal alignment.
@@
The main reason for multilingual erosion is that the data used for multimodal alignment is highly imbalanced at the language level. Due to the dominance of English-centric data, while the model aligns visual and textual tokens well in English, it performs poorly in other languages. Hence, it is crucial to align visual and textual tokens compatibly at the language level. We hypothesize that multilingual erosion may arise from the lack of alignment between visual tokens and textual tokens in other languages. From the perspective of pre-trained datasets, OpenAI-CLIP [51] is trained on the large-scale image-text pairs through contrastive learning, with the text corpus being mostly in English, potentially biasing image encoding towards an English semantic space. As shown in Figure 1, we train two separate models using the same data: one with OpenAI-CLIP vision encoder and the other with Chinese-CLIP [62] vision encoder. Interestingly, the model equipped with OpenAI-CLIP struggles to generate suitable outputs according to Chinese inputs while the other model with Chinese-CLIP can not only understand the queries but also generate appropriate outputs in Chinese. Furthermore, we observed an improvement in performance, from 66.4 to 68.3, on the MMBench-CN [39] dataset when using Chinese-CLIP. Therefore, the challenge arises: how to use English-centric multilingual image-text data to bridge the modality gap while enhancing the MLLM’s multilingual capabilities.
@@
Due to the scarcity of non-English multimodal data (e.g., lack of large-scale, high-quality image-text data), we require almost the same amount of image-text data as LLaVA to enhance the model’s multilingual capabilities. Moreover, motivated by preliminary experiments, it is necessary to condition the visual tokens on diverse language inputs. In this paper, we introduce Parrot, a novel method that utilizes textual guidance to drive visual token alignment at the language level and converts visual tokens into language-specific embeddings using a Mixture-of-Experts (MoE) module [25; 56]. Specifically, we first calculate the cross-attention between the class token of visual features extracted by the vision encoder and the text embeddings derived from word token embeddings. The result is then passed through the router of MoE to obtain the activated probability distribution of each language expert. Subsequently, demanding the input language, the English-biased visual tokens are converted into language-specific embeddings using the selected experts. This enables Parrot not only to enhance its multilingual capabilities but also to bridge the multimodal gap effectively.
@@
To address the scarcity of current multilingual benchmarks, we introduce a new benchmark encompassing six languages: English, Chinese, Portuguese, Arabic, Turkish, and Russian. This includes an extension of the MMBench-DEV dataset to these six languages and a Massive Multilingual Multimodal Benchmark (MMMB) featuring 2,000 evaluation questions per language, totaling 12,000 questions. We compare several open-source multimodal methods and evaluate some private models. Extensive experiments validate the Parrot’s state-of-the-art performance across two multilingual benchmarks. Specifically in Turkish and Arabic, our method even outperforms LLaVA-NeXT [37] by more than 10 percentage points in both benchmarks. Additionally, we evaluate our model across a broad range of multimodal benchmarks (e.g., MME [22], ScienceQA-IMG [41], and SEED-Bench-IMG [28]), demonstrating its competitive performance in diversity tasks.
2 Related Work
Multimodal Large Language Models. The domain of MLLMs has witnessed significant advances, particularly in the enhancement of visual and language processing. Current MLLM is usually a combination of visual encoders [51; 57; 21; 70; 48; 68], LLMs, and fusion modules. Innovations like Flamingo [2] and OpenFlamingo [4] have advanced visual representation by integrating a Perceiver Resampler with vision encoders. BLIP-2 [31] and InstructBLIP [17] employ Q-Former to connect the frozen LLM and vision encoder. InternVL [14] trains huge ViT and QFormer to integrate visual modalities through a multi-stage training method. MiniGPT4 [75] leverages both a Q-Former and a linear projector to bridge the gap between the vision module and LLM. Furthermore, LLaVA [38] adopts a simple MLP projector to promote the alignment between the LLM and vision encoder. mPLUG-Owl [64] introduces an approach that begins to finetune the vision encoder and align visual features, followed by tuning the LLM using LoRA [23]. Qwen-VL [6] improves visual module resolution to 448, aiming to refine the model’s visual processing capabilities. Fuyu-8B [7] directly projects image patches before integration with LLM. MM1 [43] has conducted ablative studies on connector design choices, revealing that the modality adapter type is less critical than the number of visual tokens and the resolution. MiniGemini [34] utilizes high-resolution visual tokens and high-quality data to narrow the performance gap with GPT-4 and Gemini. With the rapid advancements in open-source models, proprietary models such as GPT-4V/4o [46; 47], Gemini [58; 54], Qwen-VL-Plus/MAX [6], and Claude3 [3] have achieved outstanding results in evaluations and practical applications. In this work, owing to the simplicity of the LLaVA architecture, we adopt a framework similar to LLaVA to design our model.
Multilingual Multimodal Models. Recent years have witnessed rapid progress in the expansion of multimodal models to include a wider variety of languages. M3[44] leverages English as a pivot and alternates between English-only vision-language pre-training and multilingual masked language modeling. In contrast, UC2 [74] translates English captions into various languages and uses images as the anchor. mCLIP [10] enhances the CLIP model by aligning it with a multilingual text encoder through knowledge distillation. Thanks to the expansion of the overall capabilities of large language models [1; 5; 26; 67], their multilingual capacities have significantly improved. Integrating multilingual LLMs with visual abilities has increasingly become a research focus. In the domain of LLMs, PaLI [13] develops a 17B multilingual language-image model that spans over 100 languages. Ying-VLM [33] discovers that instruction tuning in English can extend its applicability to other languages. Ziya-Visual [40] illustrates the translation of English image-text datasets into Chinese, using in-context learning for instruction-response generation. VisCPM [24] introduces a training paradigm that fine-tunes the MLLM in a quasi-zero-shot manner based on a strong multilingual large language model. Despite these advancements, they are primarily confined to two languages or rely on the massive translated corpus. On the other hand, there is no suitable multilingual benchmark for MLLMs to evaluate the performance of multiple languages. There are also some multilingual research studies in other domains, such as multilingual machine translation [73; 49; 50; 69].
@@ 3 MMMB: A Massive Multilingual Multimodal Benchmark
In this section, we first discuss the limitations of existing benchmarks and then present the characteristics that an ideal multilingual benchmark should possess. Furthermore, we design and construct a new benchmark and provide its corresponding evaluation strategy.
@@ 3.1 Limitations of Existing Benchmarks
There are several existing multilingual benchmarks (e.g., Multi30K [20], M3Exam [71], MMBench [39], and LLaVA-Bench [38; 24]) for MLLMs, but they have some limitations: 1) Outdated Benchmarks. Multi30k is designed for image-text retrieval tasks, and the performance has nearly reached the upper bound due to the relatively easy problems. 2) Non-Standardized Evaluations. Other benchmarks, like LLaVA-Bench, rely on evaluations using GPT-4. Dependence on GPT-4 as a de facto “Ground Truth” may hinder reproducibility. Meanwhile, since LLaVA uses a deprecated version (GPT-4-0314), using other different versions could result in unfair comparisons. On the other hand, because M3Exam does not offer consistent test samples across different languages, it cannot ensure whether poor performance is due to the problem’s difficulty or the model’s lack of multilingual capabilities. 3) Limited Languages. MMBench and LLaVA-Bench are limited in English and Chinese, which can not measure the multilingual capabilities across a broad spectrum.
@@
Figure 2: Some bad cases for multilingual benchmark. Left: code reasoning is strongly related to English. Middle: logical reasoning is too challenging. Right: lack relevance between image and text.
3.2 Construction of the Multilingual Benchmark
@@
To more suitably evaluate the multilingual capabilities of MLLMs, an ideal benchmark should exhibit the following characteristics:
@@
1) Languages with Significant Differences. It should cover a diverse array of language families, selecting languages that are as distinct and non-repetitive as possible. This ensures a broad assessment of MLLMs’ ability to adapt across linguistic variances.
@@
2) Problems with Medium Level of Difficulty. The problems should not be too difficult (e.g., logical reasoning) because the aim is to assess the multilingual understanding, processing, and generating capabilities of MLLMs, not logical reasoning skills.
@@
3) Tasks with Multilingual and Multimodal. As shown in Figure 2, data within datasets should not be strongly related to English (e.g., code reasoning). It cannot be inherently transformed into multiple languages since they are composed of English words. Moreover, images should be an indispensable part when MLLMs answer the question. For instance, if given a map of the United States and asked to identify its capital, MLLMs only require the text-only ability to answer this question. Therefore, it is essential that questions highlight a significant correlation between images and texts.
@@
4) Content Consistency across Languages. The goal of this benchmark is to evaluate the multilingual capabilities of MLLMs, and we aim to show the discrepancies across different languages fairly. For example, if English questions mainly focus on addition within one hundred while Chinese questions mainly concern calculus computation, it becomes difficult to ascertain whether poor performance in Chinese arises from the complexity of the problem or the limited multilingual capabilities of MLLMs. Hence, it is crucial to ensure content consistency across languages for a fair comparison.
@@
We selected six languages for inclusion: English (en), Chinese (zh), Portuguese (pt), Arabic (ar), Turkish (tr), and Russian (ru). These languages represent a diverse range of linguistic families, and we list the detailed information and some multilingual cases in Figure 3. In terms of dataset requirements and consistency, our benchmark incorporates datasets in two main respects: 1) Since MMBench [39] officially includes English and Chinese versions, we extend it to the other four languages. 2) For the massive multilingual multimodal benchmark, denoted as MMMB, we select and clean the suitable data from ScienceQA [41], MME [22], and SEED-Bench [28] datasets with established guidelines. These datasets are then processed into a Visual Question Answering (VQA) format, resulting in a total of 12,000 samples across all six languages.
@@ 3.3 Evaluation Strategy
Since random guessing can lead to $\sim$25Top-1 accuracy for 4-choice questions, potentially reducing the discernible performance differences between various MLLMs. Additionally, MLLMs may prefer to predict a certain choice among all given choices [39], which further amplifies the bias in evaluation. To address these issues, we implement a circular validation strategy inspired by MMBench. Specifically, MMMB is adapted to the format of Yes/No questions, where each image is paired with two questions, demanding ‘Yes’ and ‘No’ answers, respectively. As shown in Figure 8, an answer is considered accurate only if both questions are answered correctly; failing either results in marking the entire instance as incorrect. This strategy ensures a more rigorous evaluation of MLLMs, reducing the likelihood of random guessing and promoting more validated comparisons across different models.
Figure 3: Overview of MMMB. It incorporates 6 languages, 15 categories, and 12,000 questions.
@@ 4.1 Preliminaries: Visual Instruction Tuning
A representative work in MLLMs is LLaVA [38], which introduces a simple yet effective method for achieving alignment between the vision encoder and the pre-trained LLM. Specifically, for a given input image ${\bf X}_{\texttt{v}}$, LLaVA utilizes the pre-trained CLIP vision encoder ViT-L/14 [51] to extract the visual features ${\bf Z}_{\texttt{v}}=g({\bf X}_{\texttt{v}})$. It then employs Vicuna [15] as the LLM to obtain the textual embeddings ${\bf H}_{t}$. To align the vision encoder with the LLM, a projector in the form of a multi-layer perceptron (MLP) denoted as ${\bf W}$ is learned. This projector converts ${\bf Z}_{\texttt{v}}$ into language embedding tokens ${\bf H}_{\texttt{v}}$, effectively facilitating the integration of multimodal information within the LLM’s framework.

### v3

# Hai-Long Sun    Da-Wei Zhou    Yang Li    Shiyin Lu    Chao Yi    Qing-Guo Chen    Zhao Xu    Weihua Luo    Kaifu Zhang    De-Chuan Zhan    Han-Jia Ye

@@ Abstract
The rapid development of Multimodal Large Language Models (MLLMs), such as GPT-4o, marks a significant step toward artificial general intelligence. Existing methods typically align vision encoders with LLMs via supervised fine-tuning (SFT), but this often deteriorates their ability to handle multiple languages as training progresses. We empirically observe that imbalanced SFT datasets, largely English-centric, degrade performance on non-English languages due to the failure in multilingual token alignment. To address this, we propose Parrot, a novel approach that leverages textual guidance for visual token alignment at the language level. Parrot conditions visual tokens on diverse language inputs and uses Mixture-of-Experts (MoE) to align multilingual tokens. By computing cross-attention between initial visual features and textual embeddings, we select the most relevant experts, converting visual tokens into language-specific representations. Additionally, we introduce the Massive Multilingual Multimodal Benchmark (MMMB), a new benchmark comprising 6 languages, 15 categories, and 12,000 questions, to assess multilingual capabilities. Parrot achieves state-of-the-art performance on both the multilingual benchmarks and a wide range of multimodal tasks. Code and dataset are available at: https://github.com/AIDC-AI/Parrot.
Machine Learning, ICML
@@ 1 Introduction
The rapid development of Large Language Models (LLMs), such as GPT-4 (Radford et al., 2018; Brown et al., 2020; OpenAI, 2023a, 2024), has gained significant attention. However, LLMs are limited to processing only text-only data. The integration of visual modalities has endowed LLMs with multimodal capabilities (Wang et al., 2024; Liu et al., 2024), driving the emergence of Multimodal Large Language Models (MLLMs). These models combine pre-trained LLMs with vision encoders, bridging the modality gap by aligning visual features with language embeddings. Current research predominantly uses either Q-Former (Li et al., 2023b) or MLP projector (Liu et al., 2023b) to align vision encoders with LLMs, enabling models to process multimodal.
Multilingual capability is a crucial aspect of MLLMs, enabling them to generate responses in the same language as the input, accommodating linguistic diversity. This feature is vital for ensuring equitable access to technology across different regions and cultures (Chen et al., 2022; Hu et al., 2023). Many LLMs (Grattafiori et al., 2024; Yang et al., 2024; OpenAI, 2023b) exhibit multilingual capabilities, generating diverse language responses based on prompts. However, after multimodal alignment training, MLLMs often lose their ability to effectively understand, process, or generate non-English languages. This phenomenon, which we refer to as multilingual erosion, results in models like LLaVA (Liu et al., 2023b), which tend to respond predominantly in English, even when the input is in another language. Therefore, it is essential to address multilingual erosion for improving MLLM’s multilingual abilities.
The primary cause of multilingual erosion is the imbalanced nature of multimodal alignment data, which is overwhelmingly English-centric. While models align visual and textual tokens well in English, their performance in other languages is suboptimal. Through empirical analysis, we observe that multilingual erosion stems from the misalignment between visual tokens and textual tokens in non-English languages. For example, in our ablation study using OpenAI-CLIP and Chinese-CLIP (Yang et al., 2022), we find that a model using OpenAI-CLIP struggles with Chinese inputs, while the Chinese-CLIP-equipped model effectively understands and generates Chinese responses. As shown in Figure 6, the t-SNE visualizations further demonstrate that the visual features of Chinese-CLIP-based LLaVA are more closely aligned with the Chinese prompts. Therefore, an intuitive question is: how to transform the visual features into language-specific embeddings to enhance the MLLM’s multilingual capabilities.
@@
Due to the scarcity of non-English multimodal data (e.g., the lack of large-scale, high-quality image-text datasets), it is necessary to use as little multilingual data as possible to enhance the model’s multilingual capabilities. To this end, we propose a novel method, Parrot, which uses textual guidance to align visual tokens at the language level. Parrot leverages a Mixture-of-Experts (MoE) module (Jacobs et al., 1991) to convert visual tokens into language-specific embeddings. Specifically, we first calculate the cross-attention between the class token of visual features and the text embeddings. The resulting features are then passed through the MoE router to activate a probability distribution for each language expert. Based on the input language, visual tokens biased towards English are transformed into language-specific embeddings using the appropriate expert. This approach not only enhances the multilingual capabilities of the MLLM but also bridges the multimodal gap effectively.
@@
To address the scarcity of current multilingual benchmarks, we introduce a new benchmark encompassing six languages: English, Chinese, Portuguese, Arabic, Turkish, and Russian. This includes an extension of the MMBench dataset to six languages and a Massive Multilingual Multimodal Benchmark (MMMB) featuring 2,000 evaluation questions per language, totaling 12,000 questions. Through a semi-automatic construction process, we mitigate the potential erros and noise. Extensive experiments validate the Parrot’s state-of-the-art performance across two multilingual benchmarks, surpassing Qwen2-VL and LLaVA-OneVision in multiple languages. Additionally, we evaluate our model across a broad range of multimodal benchmarks (e.g., MME (Fu et al., 2023) and ScienceQA (Lu et al., 2022), and SEED-Bench (Li et al., 2024a)), demonstrating its competitive performance in diverse tasks.
@@ 2 MMMB: A Massive Multilingual Multimodal Benchmark
In this section, we first outline the limitations of existing benchmarks and identify the key characteristics an ideal multilingual benchmark should exhibit. We then provide a detailed explanation of how to construct a new benchmark.
@@ 2.1 Limitations of Existing Benchmarks
There are several existing multilingual benchmarks (e.g., Multi30K (Elliott et al., 2016), M3Exam (Zhang et al., 2024b), MMBench (Liu et al., 2023c), and LLaVA-Bench (Liu et al., 2023b; Hu et al., 2023)) for MLLMs, but they have notable limitations: 1) Outdated Benchmarks. Multi30k is designed for image-text retrieval tasks, and its performance has nearly reached its upper bound due to relatively simple problems. 2) Non-Standardized Evaluations. Benchmarks like LLaVA-Bench rely on evaluations using GPT-4. This dependence on GPT-4 as a de facto “Ground Truth” may hinder reproducibility. Additionally, LLaVA uses a deprecated version (GPT-4-0314), which introduces potential inconsistencies if other versions are used, leading to unfair comparisons. Moreover, M3Exam lacks consistent test samples across different languages, making it difficult to determine whether poor performance results from the difficulty of the problem or the model’s limited multilingual capabilities. 3) Limited Languages. MMBench and LLaVA-Bench are restricted to English and Chinese, limiting their ability to assess multilingual capabilities across a broader range of languages.
@@
Figure 2: Some bad cases for the existing multilingual benchmark. Left: code reasoning is strongly related to English. Middle: logical reasoning is too challenging. Right: lack relevance between image and text.
2.2 The Key Characteristics of an Ideal Benchmark
@@
To more accurately evaluate the multilingual capabilities of MLLMs, an ideal benchmark should exhibit the following characteristics:
@@
1) Languages with Significant Differences. The benchmark should cover a diverse range of language families, selecting languages that are distinct and non-repetitive. This ensures a comprehensive assessment of MLLMs’ ability to adapt across linguistic variations.
@@
2) Problems with Medium Level of Difficulty. The problems should not be overly challenging (e.g., logical reasoning), as the primary goal is to assess the multilingual understanding, processing, and generating capabilities of MLLMs, rather than their reasoning abilities.
@@
3) Tasks with Multilingual and Multimodal. As shown in Figure 2, datasets should not be overly reliant on English (e.g., code reasoning). These tasks should not inherently translate into multiple languages if they are composed mainly of English words. Moreover, images should be an indispensable part when MLLMs answer the question. For instance, when shown a map of the United States and asked to identify its capital, relying solely on text-based abilities would be insufficient. Therefore, it is crucial that questions exhibit a significant interplay between images and text.
@@
4) Content Consistency across Languages. To fairly assess the multilingual capabilities of MLLMs, content across languages must remain consistent. For instance, if English questions focus primarily on addition within one hundred, while Chinese questions emphasize calculus computation, it would be difficult to determine whether poor performance in Chinese stems from the complexity of the problem or from the model’s limited multilingual capabilities. Ensuring content consistency across languages is thus essential for a fair and accurate comparison.
2.3 Construction Pipeline
Figure 3: The calibration process for constructing a multilingual benchmark consists of two stages: translation and calibration.
@@ 2.3 Construction Pipeline
Following the above criteria, we select six languages for inclusion: English (en), Chinese (zh), Portuguese (pt), Arabic (ar), Turkish (tr), and Russian (ru). These languages represent a diverse range of linguistic families. Detailed information and multilingual examples are provided in Figure 4. Regarding dataset requirements and consistency, our benchmark is constructed with two key considerations: 1) Since MMBench (Liu et al., 2023c) officially includes English and Chinese versions, we extend it to the other four languages. 2) For the creation of the new massive multilingual multimodal benchmark, we select and curate relevant data from the ScienceQA (Lu et al., 2022), MME (Fu et al., 2023), and SEED-Bench (Li et al., 2024a) datasets, adhering to established guidelines. These datasets are then transformed into a Visual Question Answering (VQA) format, resulting in a total of 12,000 samples across all six languages.
To mitigate potential errors and noise in the data acquisition process, we employ the following strategies to enhance the quality of our translations, as illustrated in Figure 3. First, we use GPT-4 to translate the original problem into the target language. Next, the initial translation is re-entered into for a re-check and refinement. This step helps identify and correct any immediate inconsistencies or inaccuracies. For manual calibration, we engage two groups of professional translators for each language involved in the study:
  • First Group for Refinement. This group consists of three language experts who independently review and refine the translations generated by GPT-4. This results in three distinct versions of each translation.
  • Second Group for Voting. The second group evaluates these refined translations and selects the most accurate one, ensuring it captures the intended meaning and nuances of the original text.
Figure 4: The overview of MMMB benchmark. It incorporates 6 languages, 15 categories, and 12,000 questions.
This calibration process significantly improves data quality by reducing errors and ensuring that translations are contextually accurate across languages. As a result, our benchmark achieves higher linguistic precision and cultural relevance, which we believe enhances the robustness of our research findings. Future versions will include additional details to further enhance readability and completeness.
@@ 2.4 Evaluation Strategy
Since random guessing can lead to $\sim$25Top-1 accuracy for 4-choice questions, it may reduce the discernible performance differences between various MLLMs. Additionally, MLLMs may have a tendency to favor a particular choice among the options (Liu et al., 2023c), further amplifying evaluation bias. To address these issues, we implement a circular validation strategy inspired by MMBench. Specifically, MMMB is adapted to the Yes/No question format, where each image is paired with two questions, requiring ‘Yes’ and ‘No’ answers, respectively. As shown in Figure 10 in Appendix, an answer is considered accurate only if both questions are answered correctly; failing either results in marking the entire instance as incorrect. This strategy ensures a more rigorous evaluation of MLLMs, reducing the impact of random guessing and promoting more robust comparisons across different models.
@@ 3.1 Preliminaries: Visual Instruction Tuning
A representative work in MLLMs is LLaVA (Liu et al., 2023b), which introduces a simple yet effective method for aligning the vision encoder and the pre-trained LLM. Specifically, for a given input image ${\bf X}_{\texttt{v}}$, LLaVA uses the pre-trained CLIP vision encoder ViT-L/14 (Radford et al., 2021) to extract visual features ${\bf Z}_{\texttt{v}}=g({\bf X}_{\texttt{v}})$. It then employs Vicuna (Chiang et al., 2023) as the LLM to generate textual embeddings ${\bf H}_{t}$. To align the vision encoder with the LLM, a projector, implemented as a multi-layer perceptron (MLP) denoted by ${\bf W}$. This projector converts ${\bf Z}_{\texttt{v}}$ into language embedding tokens ${\bf H}_{\texttt{v}}$, effectively enabling the integration of multimodal information within the LLM’s framework.